# Gradient Descent with Large Step Sizes: Chaos and Fractal Convergence Region

### v2

@@ 3 Simplified Matrix Factorization
In this section, we study gradient descent in the special case of problem (1) where $d_{y}=1$, i.e., factorizing a scalar $y$ as the inner product of two vectors as $u^{\top}v$. This and similar scalar factorization settings have served as canonical models for understanding large-step-size dynamics (Lewkowycz et al.,, 2020; Wang et al.,, 2022; Kreisler et al.,, 2023; Ahn et al.,, 2023; Zhu et al.,, 2023). Compared to these works, our analysis extends to critical step sizes, rather than restricting to bounded step sizes, and characterizes the chaos in gradient descent. Proofs for results in this section are in Appendix E.
@@ 3.1 Chaos at Large Step Size
where $u,v\in\mathbb{R}^{d}$, $d\geq 1$ and $y\in\mathbb{R}$. Problem (2) retains several key complexities of the general problem (1), including non-convexity, high-dimensionality, non-Lipschitz gradients, and an unbounded set of global minimizers $\mathcal{M}=\left\{u^{\top}v=y\right\}$.
@@
Theorem 1 provides a tight convergence condition for gradient descent in problem (2). The critical step size (3) consists of two components: The first term comes from the fact that, when $\eta|y|>1$, all global minimizers become unstable and only attract a measure-zero set. The second term is associated with the convergence region. Notice that the critical step size depends smoothly on the initialization and that for $\eta<1/|y|$ the convergence region is equal almost everywhere to an ellipsoid $\mathcal{D}_{\eta}^{\prime}$ in $\mathbb{R}^{2d}$ (see right panel of Figure 1). This result improves the sufficient condition for convergence obtained by Wang et al., (2022), $\eta<\eta^{*}_{1}(\bar{u},\bar{v})=\min\big\{\frac{1}{3|y|},\frac{4}{\|\bar{u}\|^{2}+\|\bar{v}\|^{2}+4|y|}\big\}$. For $y\neq 0$, their threshold $\eta^{*}_{1}$ is strictly smaller than the critical step size $\eta^{*}$ in (3).
@@
Theorem 1 provides a precise description of chaos in gradient descent: it has sensitive dependence on the initialization. Note that in problem (2), the squared norm of the parameter coincides with the loss sharpness $\lambda_{\max}(\nabla^{2}L)$ at global minimizers (see Appendix E.1.2). Hence, Theorem 1 shows that at critical step size, infinitesimal perturbations of the initialization can send the trajectory to a minimizer with arbitrarily large norm, sharpness or imbalance, or to a saddle. This is a hallmark of unpredictability: it is impossible to reduce the error in the prediction of the converging point by improving the precision in the specification of the initialization. We remark that this form of strong reachability from an arbitrarily small range of initial values is familiar in chaos theory, for examples, the Julia sets in complex system and the Wada basin boundaries in flows and systems defined by diffeomorphisms (see, e.g., Devaney and Eckmann,, 1987; Nusse and Yorke,, 1996; Aguirre et al.,, 2001). However, these classical frameworks depend on properties not satisfied by our setting, for instance, complex differentiability or invertibility of the system map, and thus do not apply here.
@@
Theorem 1 quantitatively measures chaos in gradient descent: the topological entropy is positive and is at least $\log 3$. This implies that, roughly, the number of distinct gradient descent trajectories of length $N$ grows at a rate of $3^{N}$ (further interpretation for topological entropy is provided in Appendix B.1). Beyond entropy, another aspect of trajectory complexity is shown: gradient descent admits periodic orbits of any positive integer period. This property is closely related to Li-Yorke chaos and was shown for gradient descent in univariate loss functions (Kong and Tao,, 2020; Chen et al., 2024b, ). Note, the parameter space of problem (2) has dimension $2d$ with $d\geq 1$.
@@
While Theorem 1 identifies the convergence region up to a measure-zero set, we now provide a complete description of the region. In Appendix E.1.1, we showed that every minimizer with squared norm larger than $2/\eta$ is Lyapunov-unstable and that, the basins of attraction of unstable minimizers and of the saddle have measure zero. Note, by Theorem 1, these measure-zero basins intersect $\partial\mathcal{D}_{\eta}^{\prime}$ in arbitrarily small neighborhoods and exhibit fractal structure (see left panel of Figure 2). The full convergence region $\mathcal{D}_{\eta}$ is then the union of the smooth domain $\mathcal{D}_{\eta}^{\prime}$ and the basins of unstable minimizers, excluding the basin of with the saddle $(\boldsymbol{0},\boldsymbol{0})$ when $y\neq 0$. We remark that, due to their zero measure, the basins of unstable minimizers and of the saddle cannot be easily detected by simulating training and were not been identified in prior work (Zhu et al.,, 2023). We visualize them with specialized techniques, with details provided in Appendix H.
@@
Figure 2: Left: Gradient descent applied to $L(x,y)=(xy-1)^{2}$ with $(x,y)\in\mathbb{R}^{2}$. Blue lines and purple lines represent the basins of attraction of unstable minimizers and of the saddle $(\boldsymbol{0},\boldsymbol{0})$, respectively. Right: For the same problem, we evenly sample initial values in an neighborhood on $\partial\mathcal{D}_{\eta}^{\prime}$ (the blue square). We report the distributions of the squared norm and imbalance of the converged minimizer $(x^{*},y^{*})$, and of the number of iterations to reach a loss below $10^{-8}$.
Additional experiments are conducted for the distributional behavior of gradient descent. In the right panel of Figure 2, we evenly sampled $400^{2}$ initial values in a small neighborhood on $\partial\mathcal{D}^{\prime}_{\eta}$ and ran gradient descent for problem (2). For initial values that converge, we report the distribution of the squared norm and imbalancedness of the converged minimizers, as well as the number of iterations for convergence (reaching a preset convergence criterion). Notice that as predicted the norms are bounded by $2/\eta$. Notably, even though the initial values are drawn from a very small neighborhood, the distributions have wide supports. This implies sensitivity to initialization not only at the point level but also at the distribution level. Meanwhile, the distributions are not uniform, which means that gradient descent exhibits a form of distributional implicit bias under large step sizes. We leave a closer investigation of this interesting phenomenon to future work.
@@ 3.2 Regularization Induces Fractal Convergence Boundary
where $u,v\in\mathbb{R}^{d}$, $d\geq 1$ , $\lambda\geq 0$ and $y\in\mathbb{R}$. The added regularization makes the set of global minimizers a bounded set. In particular, for problem (4), $\mathcal{M}=\{u=\mathrm{sgn}(y)v,\|u\|_{2}^{2}=|y|-\lambda\}$ when $\lambda<|y|$ and $\mathcal{M}=\left\{(\boldsymbol{0},\boldsymbol{0})\right\}$ when $\lambda\geq|y|$. Regularization is commonly used to mitigate unbounded minimizers and to establish convergence results (Cabral et al.,, 2013; Ge et al.,, 2017; Li et al., 2019a, ). However, and rather remarkably, we will show that for the regularized problem the global dynamics of gradient descent becomes even more unpredictable than for the unregularized problem: not only is the limiting point of convergent trajectories unpredictable but also the convergence itself.
The predictability of convergence depends on the geometry of the boundary of the convergence region. Two difficulties arise in analyzing this geometry: (i) the presence of basin of the saddle; and (ii) the high-dimensionality of the convergence boundary. Specifically, we observe that the basin of attraction of the saddle point intricately penetrates $\mathcal{D}_{\eta}$ and creates topological boundaries “within” $\mathcal{D}_{\eta}$ (see Figure 4). However, such boundaries are not of interest, since they do not separate points inside $\mathcal{D}_{\eta}$ from points outside, i.e., both sides lie in $\mathcal{D}_{\eta}$. This motivates us to instead consider the boundary of $\mathcal{D}_{\eta}^{\prime\prime}=\mathcal{D}_{\eta}\cup\mathcal{S}_{\eta}$, where $\mathcal{S}_{\eta}$ is the basin of the saddle. Note, the smooth domain $\mathcal{D}_{\eta}^{\prime}$ in the unregularized problem plays an analogous role in clarifying the geometry of convergence region.
The second difficulty is the high dimensionality of the boundary $\partial\mathcal{D}_{\eta}^{\prime\prime}\subset\mathbb{R}^{2d}$. To address this, we identify and reduce the symmetry in $\partial\mathcal{D}^{\prime\prime}_{\eta}$. We introduce the map $T\colon\mathbb{R}^{2d}\to\mathbb{R}^{2}$, $T(u,v)=(u^{\top}v,\|u\|^{2}_{2}+\|v\|^{2}_{2})$. The fiber of $T$, i.e., the preimage of a point in $T(\mathbb{R}^{2d})$, generically forms a manifold diffeomorphic to $\mathbb{S}^{d-1}\times\mathbb{S}^{d-1}$ and hence has a regular shape (see Appendix D). In the following, we show that gradient descent dynamics are captured by their evolution across the fibers.
@@
Let $(u_{t},v_{t})_{t\geq 0}$ denote the gradient descent trajectory in problem (4) with $\lambda\geq 0$. Let $(z_{t},w_{t})=T(u_{t},v_{t})$. There exists a planar map $F\colon\mathbb{R}^{2}\to\mathbb{R}^{2}$ that only depends on $\eta,\lambda,y$ such that $(z_{t+1},w_{t+1})=F(z_{t},w_{t})$ holds for all $t\geq 0$. In particular, $(u_{t},v_{t})$ converges to $\mathcal{M}$ if and only if $(z_{t},w_{t})$ converges to $T(\mathcal{M})$, and it converges to $(\mathbf{0},\mathbf{0})$ if and only if $(z_{t},w_{t})$ converges to $(0,0)$.
The map $F$ describes how the gradient descent trajectory evolves across the fibers of $T$. Its formulation is given in Appendix D. By Proposition 2, all points lying in the same fiber share the same convergence behavior. Therefore, roughly, the boundary $\partial\mathcal{D}_{\eta}^{\prime\prime}$ can be constructed by attaching fibers of $T$ to the the projected boundary $T(\partial\mathcal{D}^{\prime\prime}_{\eta})$; an example will be given below. Note, as the fibers are generically smooth manifolds, any geometric complexity of $\partial\mathcal{D}_{\eta}^{\prime\prime}$ will be captured by $T(\partial D_{\eta}^{\prime\prime})$.
@@
Theorem 3 (Regularized Scalar Factorization).
Consider gradient descent with step size $\eta$ for problem (4) with $0<\lambda\leq\min\{(1/\eta)-|y|,1/(2\eta)\}$. Consider the map $T(u,v)=(u^{\top}v,\|u\|^{2}_{2}+\|v\|^{2}_{2})$. Let $\mathcal{S}_{\eta}$ be the basin of attraction of $(\boldsymbol{0},\boldsymbol{0})$, which is a saddle if $\lambda<|y|$, and let $\mathcal{D}_{\eta}^{\prime\prime}=\mathcal{D}_{\eta}\cup\mathcal{S}_{\eta}$.
@@
  • Sensitivity to Initialization: For any $\theta\in\mathcal{D}_{\eta}$, the algorithm converges either to the closest global minimizer $p^{-}(\theta)=\arg\min_{p\in\mathcal{M}}\|p-\theta\|^{2}$, or the farthest $p^{+}(\theta)=\arg\max_{p\in\mathcal{M}}\|p-\theta\|^{2}$ ($p^{+}(\theta)\neq p^{-}(\theta)$ when $\lambda<|y|$). Moreover, there exist infinitely many points on $\partial\mathcal{D}_{\eta}^{\prime\prime}$ such that for any open set $O$ containing such a point, there exist $\theta^{\prime},\theta^{\prime\prime}\in O$ such that $\mathrm{GD}^{N}(\theta^{\prime})$ converges to $p^{-}(\theta^{\prime})$ and $\mathrm{GD}^{N}(\theta^{\prime\prime})$ converges to $p^{+}(\theta^{\prime\prime})$, as $N$ tends to infinity.
@@
Figure 3: Gradient descent is applied to $L(u,v)=(uv-0.5)^{2}/2+0.1(u^{2}+v^{2})$ where $(u,v)\in\mathbb{R}^{2}$. Left: the projected convergence boundary $T(\partial\mathcal{D}_{\eta}^{\prime\prime})$ is self-similar with degree three: it is covered by three smaller copies of itself (green, red, blue). Middle: The convergence boundary $\partial\mathcal{D}_{\eta}^{\prime\prime}$ consists of four replicates of $T(\partial\mathcal{D}_{\eta}^{\prime\prime})$, separated by gray lines. The only two minimizers are shown as red triangle and blue circle. Points are colored red if they converge to the red triangle, and blue if they converge to the blue circle. Right: the box-counting dimension of $T(\partial\mathcal{D}_{\eta})$ is estimated as $1.249$.
@@
Theorem 3 characterizes the geometry of the convergence boundary for problem (4). Specifically, $\partial\mathcal{D}_{\eta}^{\prime\prime}$ can be understood as a fiber-bundle-like object [^2] It is not a rigorous fiber bundle since the fiber $T^{-1}(z,w)$ might degenerate when $(z,w)$ is a singular value.: fibers of $T$, which are generically smooth and bounded manifolds, are attached to a self-similar set $T(\partial\mathcal{D}_{\eta}^{\prime\prime})$ with degree three (see the left panel of Figure 3). The case of $d=1$, i.e., $u,v\in\mathbb{R}$, is illustrated in the middle panel of Figure 3: in this case, the fibers are generically a four-point set and hence, $\partial\mathcal{D}_{\eta}^{\prime\prime}$ simply consists of four copies of $T(\partial\mathcal{D}_{\eta}^{\prime\prime})$. To quantify the fractality, we estimate the box-counting dimension of $T(\partial\mathcal{D}_{\eta}^{\prime\prime})$, which turns out to be $1.249$ (see the right panel of Figure 3). This non-integer dimension implies that $T(\partial\mathcal{D}_{\eta}^{\prime\prime})$ is essentially more complex than any smooth curve, yet fails to occupy any planar area. The fractal boundary marks unpredictability in convergence: when the initial point lies near the boundary, it is practically impossible to determine whether it is inside or outside the convergence region. This unpredictability is also quantified by the box-counting dimension, as explained in Appendix B.2.
@@
Theorem 3 shows that, up to a measure-zero set, the convergence region has an unbounded interior. This sharply contrasts with the convergence region in the unregularized case, which coincides almost everywhere with a bounded domain. By Theorem 3, gradient descent converges provided that the $u^{\top}v$ decays exponentially fast as a function of the squared norm $\left\lVert u\right\rVert^{2}_{2}+\left\lVert v\right\rVert^{2}_{2}$. Geometrically, this creates an outward spike in the convergence region. Then, the self-similarity replicates this spike infinitely many times and at multiple scales, giving rise to the spiky convergence boundary observed in Figure 3. Although Theorem 3 shows the unboundedness only for the case $y=0$, we observe qualitatively the same convergence region for general targets (for an example, see left panel in Figure 1).
@@
Fractal basin boundaries have been extensively studied in dynamical systems (see, e.g., Grebogi et al., 1983b, ; McDonald et al.,, 1985). However, these classical approaches are either largely case-specific or rely on properties that do not hold in our settings, for instance, invertibility of the system map. A more detailed discussion is in Appendix A. To our knowledge, our result provides the first rigorous characterization of a fractal convergence region in the context of machine learning optimization.
@@
Theorem 3 also shows that, although regularization eliminates unbounded global minimizers, the selected minimizer remains unpredictable. Specifically, the algorithm always selects either the minimal distance solution or the maximal distance solution over the set of global minimizers. However, this selection becomes unpredictable near the convergence boundary, as both choices can occur in arbitrarily small neighborhood (see middle panel of Figure 3). This stands is contrast with gradient descent under small step sizes, which typically is biased to the minimal distance solution (see, e.g., Gunasekar et al.,, 2018; Boursier et al.,, 2022). In fact, in Theorem 4 below, we show that this bias appears when the step size is small enough.
In the following result, we show that both the convergence and the converged minimizer are predictable when step size is sufficiently small.

### v3

@@ 3 Simplified Matrix Factorization
In this section, we study gradient descent in the special case of problem (1) where $d_{y}=1$, i.e., factorizing a scalar $y$ as the inner product of two vectors as $u^{\top}v$. This and similar scalar factorization settings have served as canonical models for understanding large-step-size dynamics (Lewkowycz et al., 2020; Wang et al., 2022; Kreisler et al., 2023; Ahn et al., 2023; Zhu et al., 2023; Kalra et al., 2025). Compared to these, our analysis extends to the full spectrum of step sizes, rather than restricting to bounded step sizes. Proofs for results in this section are in Appendix E.
@@ 3.1 Chaos at Large Step Size
where $u,v\in\mathbb{R}^{d}$, $d\geq 1$ and $y\in\mathbb{R}$. Problem (2) retains several key complexities of the general problem (1), including non-convexity, high-dimensionality, non-Lipschitz gradients, and an unbounded set of global minimizers $\mathcal{M}=\left\{\theta\colon u^{\top}v=y\right\}$.
@@
Theorem 1 provides a necessary and sufficient convergence condition for gradient descent in problem (2). The critical step size (3) consists of two terms. The first term arises because: when $\eta|y|>1$, all global minimizers become unstable and only attract a measure-zero set. The second term characterizes the full convergence region, which is equal almost everywhere to an ellipsoid $\mathcal{D}_{\eta}^{\prime}$ (see right panel of Figure 1). Note that this convergence result is global: it holds for all step sizes $\eta$ and initializations $\theta_{0}$, except for a null set in the $(\eta,\theta_{0})$-space. In contrast, the previous work of Wang et al. (2022) only showed convergence for step sizes in the range $\eta<1/(3|y|)$, and initializations in a strict subset of $\mathcal{D}_{\eta}^{\prime}$ (see Figure 7). A comparison between the proof techniques used for Theorem 1 and those used by Wang et al. (2022) is provided in Appendix E.1.2.
@@
By Theorem 1, gradient descent in problem (2) exhibits a strong form of sensitivity to initial condition. Note that in problem (2), the squared norm of the parameter coincides with the loss sharpness $\lambda_{\max}(\nabla^{2}L)$ at global minimizers (see Appendix E.1.2). Hence, Theorem 1 shows that at critical step size, infinitesimal perturbations of the initialization can send the trajectory to a minimizer with arbitrarily large norm, sharpness or imbalance, or to a saddle. This is a hallmark of unpredictability: it is impossible to reduce the error in the prediction of the converging point by improving the precision in the specification of the initialization. We remark that this form of reachability from an arbitrarily small range of initial values is familiar in chaos theory, such as the Julia sets in complex systems and the Wada basin boundaries (see, e.g., Devaney and Eckmann, 1987; Nusse and Yorke, 1996; Aguirre et al., 2001). However, these classical frameworks depend on properties not satisfied by our setting, such as complex differentiability or invertibility of the system map, and thus do not apply here.
@@
Theorem 1 also quantitatively measures chaos in gradient descent in problem (2): the topological entropy is positive and is at least $\log 3$. Roughly speaking, the number of distinct gradient descent trajectories of length $N$ grows at a rate of $3^{N}$ (further interpretation for topological entropy is provided in Appendix B.1). Beyond entropy, another aspect of trajectory complexity is shown: gradient descent admits periodic orbits of any positive integer period. This is closely related to Li-Yorke chaos and was shown for gradient descent in univariate loss functions (Kong and Tao, 2020; Chen et al., 2024b).
@@
While Theorem 1 identifies the convergence region up to a measure-zero set, we now provide a complete description of the region. In Appendix E.1.1, we show that every minimizer with squared norm larger than $2/\eta$ is Lyapunov-unstable and that the basins of attraction of unstable minimizers and of the saddle have measure zero. By Theorem 1, these measure-zero basins intersect $\partial\mathcal{D}_{\eta}^{\prime}$ in arbitrarily small neighborhoods and exhibit fractal structure (see left panel of Figure 2). The full convergence region $\mathcal{D}_{\eta}$ is then the union of the smooth domain $\mathcal{D}_{\eta}^{\prime}$ and the basins of unstable minimizers, excluding the basin of the saddle $(\boldsymbol{0},\boldsymbol{0})$ when $y\neq 0$.
@@
Figure 2: Gradient descent applied to $L(u,v)=(uv-1)^{2}$ with $(u,v)\in\mathbb{R}^{2}$. Left: Blue lines and purple lines represent the basins of attraction of unstable minimizers and of the saddle $(\boldsymbol{0},\boldsymbol{0})$, respectively. Right: Initializations are evenly sampled from a small set intersecting $\partial\mathcal{D}_{\eta}^{\prime}$ (the blue square). However, the distributions of the squared norm and imbalance of the converged minimizer $(u^{*},v^{*})$, and of the number of iterations to reach a loss below $10^{-8}$, have wide supports.
The next result shows that gradient descent in problem (2) exhibits chaos even after removing the measure-zero basins of attraction of unstable minimizers, i.e., it is chaotic at the distributional level.
Theorem 2.
Under the same conditions and notations as Theorem 1 and assuming $\eta|y|<1$, given arbitrary $\theta\in\partial\mathcal{D}_{\eta}^{\prime},\varepsilon>0$ and open sub-interval $E\subset[\gamma_{\min},2/\eta]$, there exists a set of positive measure $O\subset B(\theta,\varepsilon)$ such that for any $\theta^{\prime}\in O$, $\mathrm{GD}_{\eta}^{N}(\theta^{\prime})$ converges to a solution with norm in $E$.
Consider initializations randomly sampled from a distribution whose support contains an arbitrarily small neighborhood of the convergence boundary. Theorem 2 then implies that the sharpness (or norm) of the converged minimizer must have a support containing the entire interval $(\gamma_{\min},2/\eta)$ (see right panel of Figure 2). In other words, at near-critical step size and even after discarding the basins of unstable minimizers, the best prediction for the final sharpness is the entire interval $(\gamma_{\min},2/\eta)$.
@@ 3.2 Regularization Induces Fractal Convergence Boundary
where $u,v\in\mathbb{R}^{d}$, $d\geq 1$ , $\lambda\geq 0$ and $y\in\mathbb{R}$. The added regularization makes the set of global minimizers a bounded set. In particular, for problem (4), $\mathcal{M}=\{u=\mathrm{sgn}(y)v,\|u\|_{2}^{2}=|y|-\lambda\}$ when $\lambda<|y|$ and $\mathcal{M}=\left\{(\boldsymbol{0},\boldsymbol{0})\right\}$ when $\lambda\geq|y|$. Regularization is commonly used to mitigate unbounded minimizers and to establish convergence results (Cabral et al., 2013; Ge et al., 2017; Li et al., 2019a). However, and rather remarkably, we will show that for the regularized problem the global dynamics of gradient descent becomes even more unpredictable than for the unregularized problem: not only is the limiting point of convergent trajectories unpredictable but also the convergence itself.
The predictability of convergence depends on the geometry of the boundary of the convergence region. Two difficulties arise in analyzing this geometry: (i) the presence of the basin of the saddle; and (ii) the high-dimensionality of the convergence boundary. Specifically, we observe that the basin of attraction of the saddle point intricately penetrates $\mathcal{D}_{\eta}$ and creates topological boundaries “within” $\mathcal{D}_{\eta}$ (see Figure 8). However, such boundaries are not of interest, since they do not separate points inside $\mathcal{D}_{\eta}$ from points outside, i.e., both sides lie in $\mathcal{D}_{\eta}$. This motivates us to instead consider the boundary of $\mathcal{D}_{\eta}^{\prime\prime}=\mathcal{D}_{\eta}\cup\mathcal{S}_{\eta}$, where $\mathcal{S}_{\eta}$ is the basin of the saddle. Note, the smooth domain $\mathcal{D}_{\eta}^{\prime}$ in the unregularized problem plays an analogous role in clarifying the geometry of convergence region.
The second difficulty is the high dimensionality of the boundary $\partial\mathcal{D}_{\eta}^{\prime\prime}\subset\mathbb{R}^{2d}$. To address this, we identify and reduce the symmetry in $\partial\mathcal{D}^{\prime\prime}_{\eta}$. We introduce the map $T\colon\mathbb{R}^{2d}\to\mathbb{R}^{2}$, $T(u,v)=(u^{\top}v,\|u\|^{2}_{2}+\|v\|^{2}_{2})$. The fiber of $T$, i.e., the preimage of a point in $T(\mathbb{R}^{2d})$, is generically a manifold diffeomorphic to $\mathbb{S}^{d-1}\times\mathbb{S}^{d-1}$ and hence has a regular shape (see Appendix D). In the following, we show that gradient descent dynamics are captured by their evolution across these fibers.
@@
Let $(u_{t},v_{t})_{t\geq 0}$ denote the gradient descent trajectory in problem (4) with $\lambda\geq 0$. Let $(z_{t},w_{t})=T(u_{t},v_{t})$. There exists a planar map $F\colon\mathbb{R}^{2}\to\mathbb{R}^{2}$ that only depends on $\eta,\lambda,y$ such that $(z_{t+1},w_{t+1})=F(z_{t},w_{t})$ holds for all $t\geq 0$. In particular, $(u_{t},v_{t})$ converges to $\mathcal{M}$ if and only if $(z_{t},w_{t})$ converges to $T(\mathcal{M})$, and it converges to $(\boldsymbol{0},\boldsymbol{0})$ if and only if $(z_{t},w_{t})$ converges to $(0,0)$.
The map $F$ describes how the gradient descent trajectory evolves across the fibers of $T$. Its formulation is given in Appendix D. By Proposition 3, all points lying in the same fiber share the same convergence behavior. Therefore, roughly, the boundary $\partial\mathcal{D}_{\eta}^{\prime\prime}$ can be constructed by attaching fibers of $T$ to the projected boundary $T(\partial\mathcal{D}^{\prime\prime}_{\eta})$; an example will be given below. Note, as the fibers are generically smooth manifolds, any geometric complexity of $\partial\mathcal{D}_{\eta}^{\prime\prime}$ will be captured by $T(\partial D_{\eta}^{\prime\prime})$.
@@
Theorem 4.
Consider gradient descent with step size $\eta$ for problem (4) with $0<\lambda\leq\min\{(1/\eta)-|y|,1/(2\eta)\}$. Consider the map $T(u,v)=(u^{\top}v,\|u\|^{2}_{2}+\|v\|^{2}_{2})$. Let $\mathcal{S}_{\eta}$ be the basin of attraction of $(\boldsymbol{0},\boldsymbol{0})$, which is a saddle if $\lambda<|y|$, and let $\mathcal{D}_{\eta}^{\prime\prime}=\mathcal{D}_{\eta}\cup\mathcal{S}_{\eta}$. The following holds:
@@
  • Sensitivity to Initialization: For any $\theta\in\mathcal{D}_{\eta}$, the algorithm converges either to the closest global minimizer $p^{-}(\theta)=\arg\min_{p\in\mathcal{M}}\|p-\theta\|^{2}$, or the farthest $p^{+}(\theta)=\arg\max_{p\in\mathcal{M}}\|p-\theta\|^{2}$. [^2] When $\theta\in\mathcal{D}_{\eta}$, both $\min_{p\in\mathcal{M}}\|p-\theta\|^{2}$ and $\max_{p\in\mathcal{M}}\|p-\theta\|^{2}$ admit unique solutions. Moreover, $p^{+}(\theta)\neq p^{-}(\theta)$ when $\lambda<|y|$. Moreover, there exist infinitely many points on $\partial\mathcal{D}_{\eta}^{\prime\prime}$ such that for any open set $O$ containing such a point, there exist $\theta^{\prime},\theta^{\prime\prime}\in O$ such that $\mathrm{GD}_{\eta}^{N}(\theta^{\prime})$ converges to $p^{-}(\theta^{\prime})$ and $\mathrm{GD}_{\eta}^{N}(\theta^{\prime\prime})$ converges to $p^{+}(\theta^{\prime\prime})$, as $N$ tends to infinity.
@@
Figure 3: Gradient descent is applied to $L(u,v)=(uv-0.5)^{2}/2+0.1(u^{2}+v^{2})$, where $(u,v)\in\mathbb{R}^{2}$. Left: The projected convergence boundary $T(\partial\mathcal{D}_{\eta}^{\prime\prime})$ is self-similar with degree three: it is covered by three smaller copies of itself (green, red, blue). Middle: The convergence boundary $\partial\mathcal{D}_{\eta}^{\prime\prime}$ consists of four replicates of $T(\partial\mathcal{D}_{\eta}^{\prime\prime})$, separated by gray lines. The only two minimizers are shown as a red triangle and a blue circle. Points are colored red if they converge to the red triangle, and blue if they converge to the blue circle. Right: The box-counting dimension of $T(\partial\mathcal{D}_{\eta}^{\prime\prime})$ is estimated as $1.249$.
@@
Theorem 4 indicates that the convergence boundary for problem (4), $\partial\mathcal{D}_{\eta}^{\prime\prime}$, can be constructed by attaching fibers of $T$ to a self-similar set $T(\partial\mathcal{D}_{\eta}^{\prime\prime})$ with degree three. $T(\partial\mathcal{D}_{\eta}^{\prime\prime})$ is displayed in the left panel of Figure 3, and its box-counting dimension is estimated to be $1.249$, as shown in the right panel. The case of $d=1$, i.e., $u,v\in\mathbb{R}$, is shown in the middle panel of Figure 3. There, the fibers of $T$ generically consist of four discrete points. $\partial\mathcal{D}_{\eta}^{\prime\prime}$ is then constructed by attaching four points to $T(\partial\mathcal{D}_{\eta}^{\prime\prime})$. Hence, $\partial\mathcal{D}_{\eta}^{\prime\prime}$ is simply the union of four copies of $T(\partial\mathcal{D}_{\eta}^{\prime\prime})$. The fractal boundary marks unpredictability in convergence: given an initial point near the boundary, it is almost impossible to determine whether the point is inside or outside the convergence region. This unpredictability is also quantified by the box-counting dimension, as explained in Appendix B.2.
@@
Theorem 4 shows that when $y=0$, the convergence region has an unbounded interior up to a null set. This sharply contrasts with the convergence region in the unregularized case, which coincides almost everywhere with a bounded domain. By Theorem 4, gradient descent converges provided that the $u^{\top}v$ decays exponentially fast as a function of the squared norm $\left\lVert u\right\rVert^{2}_{2}+\left\lVert v\right\rVert^{2}_{2}$. Geometrically, this creates an outward spike in the convergence region. Then, the self-similarity replicates this spike infinitely many times and at multiple scales, giving rise to the spiky convergence boundary observed in Figure 3. Although Theorem 4 shows the unboundedness only for the case $y=0$, we observe qualitatively the same geometry for general targets (for an example, see left panel in Figure 1).
@@
Fractal basin boundaries have been extensively studied in dynamical systems (see, e.g., Grebogi et al., 1983b; McDonald et al., 1985). However, these classical approaches are either largely case-specific or rely on properties that do not hold in our settings, for instance, invertibility of the system map. A more detailed discussion is in Appendix A. To our knowledge, our result provides the first rigorous characterization of a fractal convergence region in the context of machine learning optimization.
@@
Theorem 4 also shows that, although regularization eliminates unbounded global minimizers, the selected minimizer remains unpredictable. Specifically, while the algorithm always selects either the minimal distance solution or the maximal distance solution, this selection becomes unpredictable near the convergence boundary, as both choices occur in arbitrarily small sets (see middle panel of Figure 3). This stands in contrast with gradient descent under small step sizes, which typically exhibits a distance-minimization bias (see, e.g., Gunasekar et al., 2018; Boursier et al., 2022). Indeed, in the following, we show that with sufficiently small step size, gradient descent in problem (4) always selects the minimal distance solution.